# Shared Control with Black Box Agents using Oracle Queries

### v1

@@ 5 Expeiments
The third use case, Combination-lock (Figure 2(c)), is designed such that the control should learn a specific sequence of actions. In this case, querying can significantly improve training since instead of training randomly until it converges, the oracle will directly lead the control to the correct sequence. Moreover, there are no hidden actions as in the previous cases, so if the right action is selected, the learned action will also be useful at runtime. Thus, in this use case, the expectation is that querying will benefit in the learning stage, and will not improve beyond that during the testing stage.
@@
Domain 2: Lunar Lander Shared Control
The Atari-based system depicted in Figure 2(d) consists of the Lunar Lander simulation in OpenAI Gym [33]. This domain is a classic rocket trajectory optimization problem that was used in previous work to evaluate shared control algorithms designed to learn from human users’ interactions [9]. The black box $\mathbb{E}$ is the operator of the lander, and the control $\mathbb{C}$ is its fuel system. The states are shared between the operator and the control, such that $S$ is the current position of the lander (x,y, and angle) and its velocity (x,y, and angular). The operator chooses one of four actions: do nothing, engage the right engine, engage the left engine, and engage the central engine. The control has two actions: inject fuel into the engines or not. The joint operation protocol of this system is the restriction protocol, similar to the protocol presented in Example 1: {dmath*} T(sC,sE,aC,¬inject) = (nothing(sC), ¬inject (sE))
@@
Setup In both domains, for each combination of a scenario, oracle type, and a heuristic, an RNN was trained for $40$ epochs. In the automaton domain, each epoch runs on a gradually increasing number of steps, from one to $19$ steps, and in the Lunar Lander domain, it was trained for $200$ steps in each epoch, which is the length of an episode in the simulation. Each configuration was run five times during training, and five more times during testing. After some manual tuning, the heuristics thresholds are set to $\beta_{ent}=0.25$ and $\beta_{util}=0.95$. The hyperparameter values of the Q-learning approach were set to be $\alpha=0.5,\gamma=0.9,\epsilon=0.05$.
@@
6 Results
Table 1: Average failure % (lower is better) across different configurations, including Oracle Type (O) being either Teacher (T) or Expert (E). Baseline runs include “No Oracle”: the RNN learns without queries, “Random”: the control chooses a random action, “Always Train”: the control always uses the Oracle during Training, “Always Train+Test”: the control always uses the Oracle both in training and in test time. Heuristic runs include “Entropy” and “Util” are the respective heuristics with queries enabled both in training and test time, and “RL Train” and “RL Train+Test” for the RL heuristic with queries enabled in training only and in train+test, respectively.
@@ 6 Results
Table 2: Failure rate, number of oracle uses, and total reward in the Lunar Lander domain with and without using heuristics. The results using the heuristics are shown in the three rightmost columns.
@@
We will first present accuracy results, then oracle usage and we will end this section with a discussion on important results. In both Table 1 and Table 2, the left columns show the failure rate of the baselines, and the right columns show the failure rate of the heuristics: Entropy, Util, and RL. For Table 1, we show two variants of the RL heuristic (“Train” and “Train+Test”), as there was a significant difference between these conditions, especially in the Strategy use case. We start by discussing the baselines. Given the teacher oracle, the two baselines “Always Train” and“Always Train+Test” improve the accuracy of the results in two out of the three example scenarios (“strategy” and “combination lock”).However, for “cases” these baselines did not achieve better performance compared to not using any query (“No Oracle”). When using expert oracle in “strategy”, the failure rate of the “always train + test” baseline increases significantly, and in “Combination Lock” the failure rate of the “Always” baselines using the expert oracle is very high ($\sim 86\%$) since the environment can accept two actions, one of which leads to a “sink” state. The Entropy heuristic performs similarly to the “Always” variants in most cases. As we will show later, it does so by querying much less, thus being more efficient. We wish to focus on one specific scenario, to discuss the reason that the Entropy heuristic did not improve accuracy. In “Combination Lock”, the failure rate with the teacher oracle is very high compared to “Always”. We checked our failure results over the five tests, and the results are on the extreme, either 00 or $98.5\%$. The reason for this dichotomy is that a failure happens when the choice at the second step of the RNN at the combination lock can either lead us to a “sink” state (See $s_{3}$ in Figure 2(c)) or to a successful continuation of the shared operation. These extreme cases are what lead to this highly variant result. Util heuristic solves the problem of choosing actions that lead to a sink-state in this use case. The teacher oracle suggests choosing ’a’,’b’, ’a’, so the agent does not fall into the sink state and learns better. Thanks to this change, the failure rate in the Util heuristic is $0.39\pm 0.88$ compared to $39.59\pm 53.77$ in the Entropy heuristic.
@@
A key result for the RL heuristic is in the “Strategy” use case, when using the “Train+Test” variant. The failure rate decreases from 52.55 $\pm$ 5.44 to 8.3 $\pm$ 17.92 with the teacher oracle. This improvement is due to the choice of the control to query when reaching state $g_{2}$ (see Figure 2(b)), which is a state that requires a single failure if we want the joint system to minimize failure.
@@
In Table 2 we show the results of using the heuristics in the Lunar Lander domain compared to the baselines (as in Table 1). The use of heuristics was beneficial in all cases. The failure rate is much lower than not using Oracle or choosing a random action and is very similar to “Always Train+Test”, where the system always uses Oracle. Additionally, in terms of the average number of queries used decreases to around 18 when using the Entropy or Util heuristics, versus 200 $\pm$ 0 for these “Always” baselines. This result means that the system uses the oracle much less and reaches very similar results to always query. Also, the reward when using the heuristics is very similar to the one they receive when always using the oracle. Next, follow with the meaning of these results. When using queries during learning time only, all heuristics show similar performance compared to the best possible strategy in Cases and Strategy (“Always” with the teacher oracle), and both the Utility and the RL heuristics exhibit close to the best possible performance in the Combination Lock scenario. When allowing queries during test time, The RL heuristic is the only one able to learn to query at the right times to benefit from querying, even though the Q-function does not continue to change during test time. Additionally, the expert oracle has the potential to actually hurt the shared control (in the Strategy use case when always querying in train+test), and hence, a good heuristic should learn not just when to query but also when not to rely on the oracle. All heuristics learned when not to use the oracle, showing similar failure rates in all scenarios.
(a) Entropy Heuristic.
(b) Utility Heuristic.
(c) RL Heuristic.
@@
Number of queries used The test in Figure 3 examines whether the RNN decreases its reliance on queries with time. The x-axis shows the number of epochs learned, and the y-axis shows the number of initiated queries. For each query policy-oracle combination, five tests were held so there are 10 lines in each graph in total. The trends are similar in all three setups, so we focus here only on the combination lock use case. Here, the use of a teacher and an expert oracle are quite distinct from one another. With the teacher, the RNN is often able to learn a policy in which a small number of queries are sufficient, while with the expert, the control keeps querying it more often.
@@
Generalization. We use the Lunar Lander use case to evaluate how the heuristics generalize to a more realistic environment. Table 2 shows that all of the trends observed in the automata use case remain consistent: the failure rate when using queries decreases significantly compared to the same runs without queries. The total number of queries is still the highest in the RL heuristic. The new interesting information from this domain is that the reward, which is an external metric for the performance of the shared system (unlike failure rate which is an intrinsic metric) is not much different between the policy without queries and with them.
7 Conclusions
@@
This paper introduces a new formalization for a shared control system with queries. It explores the new formalization by discussing and evaluating two potential oracles that can be queried, as well as three heuristic approaches for choosing when to query in such a shared system. The results show a clear trade-off between the proposed heuristics: the utility heuristic showed the best results regarding query efficiency, reducing the number of required queries, sometimes to the minimum number of queries in which the true best action cannot be learned. On the other hand, the RL heuristic can reach much lower failure rates but at the cost of a large and inconsistent number of queries. These results are consistent across two very different evaluation domains. Future work will investigate using RL techniques to make querying more efficient and consistent. One option is to replace the RNN itself with an RL algorithm that can choose between the system’s different actions and querying. Our empirical evaluation also demonstrated the impact of the oracle quality, showcasing its importance. Future work can reason about the oracle type in real time and adapt the chosen heuristic to that oracle. Lastly, another question to investigate is how to design a querying heuristic that accounts for an oracle that intentionally misleads the control.

### v2

@@ V-A1 Automata-based Shared Control
The third use case, Combination-lock (Figure 2(c)), is designed such that the control should learn a specific sequence of actions to succeed. In this case, querying can significantly improve training since instead of training the network randomly until it converges, the oracle will directly lead the control to the correct sequence. Moreover, there are no hidden actions as in the previous cases, so if the right action is selected, the learned action will also be useful at runtime. Thus, in this use case, the expectation is that querying will benefit in the learning stage, and will not improve beyond that during the testing stage.
@@
V-A2 Lunar Lander Shared Control
The Atari-based system depicted in Figure 2(d) consists of the Lunar Lander simulation in OpenAI Gym [7]. This domain is a classic rocket trajectory optimization problem that was used in previous work to evaluate shared control algorithms designed to learn from human users’ interactions [6]. The black box $\mathbb{E}$ is the operator of the lander, and the control $\mathbb{C}$ is its fuel system. The states are shared between the operator and the control, such that $S$ is the current position of the lander (x,y, and angle) and its velocity (x,y, and angular). The operator chooses one of four actions: do nothing, engage the right engine, engage the left engine, and engage the central engine. The control has two actions: inject fuel into the engines or not. The joint operation protocol of this system is the restriction protocol, similar to the protocol presented in Example 1: {dmath*} T(sC,sE,aC,inject) = (aC(sC), inject(sE)) T(sC,sE,aC,¬inject) = (nothing(sC), ¬inject (sE))
@@ V-A2 Lunar Lander Shared Control
(a) Entropy Heuristic.
(b) Utility Heuristic.
(c) RL Heuristic.
V-B Setup
In both domains, for each combination of a scenario, oracle type, and a heuristic, an RNN was trained for $40$ epochs. In the automaton domain, each epoch runs on a gradually increasing number of steps, from one to $19$ steps, and in the Lunar Lander domain, it was trained for $200$ steps in each epoch, which is the length of an episode in the simulation. Each configuration was run five times during training, and five more times during testing. After some manual tuning, the heuristics thresholds are set to $\beta_{ent}=0.25$ and $\beta_{util}=0.95$. The hyperparameter values of the Q-learning approach were set to be $\alpha=0.5,\gamma=0.9,\epsilon=0.05$.
@@ V-B Setup
VI Results
TABLE I: Average failure % (lower is better) across different configurations, including Oracle Type (O) being either Teacher (T) or Expert (E). Baseline runs include “No Oracle”: the RNN learns without queries, “Random”: the control chooses a random action, “Always Train”: the control always uses the Oracle during Training, “Always Train+Test”: the control always uses the Oracle both in training and in test time. Heuristic runs include “Entropy” and “Util” are the respective heuristics with queries enabled both in training and test time, and “RL Train” and “RL Train+Test” for the RL heuristic with queries enabled in training only and in train+test, respectively.
@@ VI Results
TABLE II: Results of failure rate, number of oracle uses, and total reward in the Lunar Lander domain with and without using heuristics. The results using the heuristics are shown in the three rightmost columns.
@@
We will first present accuracy results, then oracle usage and we will end this section with a discussion on important results. In both Table I and Table II, the left columns show the failure rate of the baselines, and the right columns show the failure rate of the heuristics: Entropy, Util, and RL. For Table I, we show two variants of the RL heuristic (“Train” and “Train+Test”), as there was a significant difference between these conditions, especially in the Strategy use case.
We first look at Table I, and discuss the baselines. Given the teacher oracle, the two baselines “Always Train” and“Always Train+Test” improve the accuracy of the results in two out of the three example scenarios (“strategy” and “combination lock”).However, for “cases” these baselines did not achieve better performance compared to not using any query (“No Oracle”). When using expert oracle in “strategy”, the failure rate of the “always train + test” baseline increases significantly, and in “Combination Lock” the failure rate of the “Always” baselines using the expert oracle is very high (approximately 86%) since the environment can accept two actions, one of which leads to a “sink” state.
So far we talked about our baselines. We now move on to discuss the heuristic. As seen in the table, the Entropy heuristic is able to reach a similar performance to the “Always” variants in most cases. As we will show later, it does so by querying much less, thus being more efficient.
@@
We wish to focus on one specific scenario, to discuss the reason that the Entropy heuristic did not improve accuracy. In “Combination Lock”, the failure rate with the teacher oracle is very high compared to “Always”. We checked our failure results over the five tests, and the results are on the extreme, getting either a 00 or a $98.5\%$ failure rate. The reason for this dichotomy is that a failure happens when the choice at the second step of the RNN at the combination lock can either leads us to a “sink” state (See $s_{3}$ in Figure 2(c)) or to a successful continuation of the shared operation. These extreme cases are what lead to this highly variant result. Util heuristic solves the problem of choosing actions that lead to a sink-state in this use case. The teacher oracle suggests choosing ’a’,’b’, ’a’, so the agent does not fall into the sink state and learns better. Thanks to this change, the failure rate in the Util heuristic is $0.39\pm 0.88$ compared to $39.59\pm 53.77$ in the Entropy heuristic.
@@
The most interesting result for the RL heuristic is in the “Strategy” use case, when using the “Train+Test” variant. The failure rate decreases from 52.55 $\pm$ 5.44 to 8.3 $\pm$ 17.92 with the teacher oracle. This improvement is due to the choice of the control to use a query when reaching state $g_{2}$ (see Figure 2(b)), which is a state that requires a single failure if we want the joint system to succeed at the rest of the trial. This policy means that the RL heuristic makes the system query at state $g_{2}$, even though it decided not to query in previous heuristics, and then often took a wrong (although myopically correct) action.
@@
In Table II we show the results of using the heuristics in the Lunar Lander domain compared to the baselines (as in Table I. From the table, it can be seen that the use of heuristics was beneficial in all cases. The failure rate is much lower than not using Oracle or choosing a random action and is very similar to “Always Train+Test”, where the system always uses Oracle. Additionally, in terms of the average number of queries used decreases to around 18 when using the Entropy or Util heuristics, versus 200 $\pm$ 0 for these “Always” baselines. This result means that the system uses the oracle much less and reaches very similar results to always query. Also, the reward when using the heuristics is very similar to the one they receive when always using the oracle. Next, follow with the meaning of these results.
VII Discussion
Querying during training. All heuristics show similar performance compared to the best possible strategy in Cases and Strategy (“Always” with the teacher oracle), and both the Utility and the RL heuristics exhibit close to the best possible performance in the Combination Lock scenario.
Querying during test time. The Entropy and Utility heuristics do not seem to leverage the ability to query during test time, even when this ability can improve the performance significantly (In Strategy and Combination Lock with the teacher oracle). The RL heuristic is able to learn to query at the right times to benefit from querying, even though the Q-function does not continue to change during test time.
Learning when not to query. The expert oracle has the potential to actually hurt the shared control (in the Strategy use case, when always querying in train+test), and hence a good heuristic should learn not just when to query, but also when not to rely on the oracle. All heuristics learned when not to use the oracle, showing similar failure rates in all scenarios.
@@ VII Discussion
Number of queries used. The test in Figure 3 examines whether the RNN decreases its reliance on queries with time. The x-axis shows the number of epochs learned so far, and the y-axis shows the number of queries the control initiated. For each query policy-oracle combination, five tests were held so there are 10 lines in each graph in total. The trends are similar in all three setups, so we focus here only on the combination lock use case. Here, the use of a teacher and an expert oracle are quite distinct from one another. With the teacher, the RNN is often able to learn a policy in which a small number of queries are sufficient, while with the expert, the control keeps querying it more often.
@@
Generalization. We use the Lunar Lander use case to evaluate how the heuristics generalize to a more realistic environment. Table II shows that all of the trends observed in the automata use case remain consistent: the failure rate when using queries decreases significantly compared to the same runs without queries. The total number of queries is still the highest in the RL heuristic. The new interesting information from this domain is that the reward, which is an external metric for the performance of the shared system (unlike failure rate which is an intrinsic metric) is not much different between the policy without queries and with them.
VIII Conclusions
This paper introduces a new formalization for a shared control system with queries. It explores the new formalization by discussing and evaluating two potential oracles that can be queried, as well as three heuristic approaches for choosing when to query in such a shared system.
@@ VIII Conclusions
The results show a clear trade-off between the proposed heuristics: the utility heuristic showed the best results regarding query efficiency, reducing the number of required queries, sometimes to the minimum number of queries in which the true best action cannot be learned. On the other hand, the RL heuristic can reach much lower failure rates but at the cost of a large and inconsistent number of queries. These results are consistent across two very different evaluation domains. Future work will investigate using RL techniques to make querying more efficient and consistent. One option is to replace the RNN itself with an RL algorithm that can choose between the system’s different actions and querying.
Our empirical evaluation also demonstrated the impact of the oracle quality, showcasing its importance. Future work can reason about the oracle type in real time and adapt the chosen heuristic to that oracle. Lastly, another question to investigate is how to design a querying heuristic that accounts for an oracle that intentionally misleads the control.